5. Consider an US firm that imports components from China, and adds value by assembling these components into a finished product in the US. Assume that the US firm always pays in US$ for the components. Suppose the cost of imported components is $5, while domestic assembly costs to finish the product are $5, and the current exchange rate is RMB5/$.
a) Case A: If the US firm was selling 100% of the finished product output in the domestic US market at $12/unit, what is the margin of the US firm as a % of revenue? (2 points)
b) Case B: If the US firm was exporting 100% of the finished product output to Eurozone markets and gets paid €12/unit in Euros, and the exchange rate is $1/€, what is the margin of the US firm as % of revenue? (2 points)
c) Case C: In case A above, if the exchange rate changed to RMB 6/$, and all the other numbers and variables remain constant, what is the margin of the US firm as % of revenue? (2 points)
d) Case D: In case B above, if the exchange rate changed to $0.9/€, and all the other numbers and
variables remain constant (RMB5/$), what is the new margin of the US firm as % of revenue. (2 points)
Answer #1

Solution :- (a)

Selling price = $12

Total cost = import + assembly = $5 + $5 = $10

So margin = $2

Now margin as a % of revenue = 2/12 = 16.67%

(b) if the firms export at 12 euro and 1 Euro = 1 dollar

Then we say selling price is 12 $ also

So the margin of revenue is same (no change) = 16.67%

As the exchange rate of dollar and euro are equal.

(C) if the Exchange rate change to RMB6/$ then cost of imported component = $4.167

$5 * RMB5/RMB6

Then total cost = $5 + $4.167 = $9.167

Then Margin = $12 - $9.167 = $2.83

Margin (%) of revenue = $2.83/$12 = 23.58%

(D)

If selling price is 12 Euro and 1$ = 0.9 Euro then

Selling price in dollar = 12 / 0.9 = $13.33

Margin = $13.33 - $10 = $3.33

The margin (%) of revenue = $3.33 / $13.33 = 25%

If there is any doubt please ask in comments.